8. In your political science class you are given a list of 10 books. You are to select 4 to read during the semester. How many different combinations of 4 books are available from the list of 10?
Answer by edjones(7569)

10C4
=10!/((10-4)!4!)
=10*9*8*7*6!/6!*4*3*2
=10*9*8*7/4*3*2
=10*3*7
=210
.
Ed